Kyla is 5 years younger than Via.Next year their ages will have a sum of 57.Find their ages.

Answer by checkley79(3341)

VIA=X
KYLE=X-5
X+1+X-5+1=57
2X=57-1+5-1
2X=60
X=60/2
X=30 YEARS OLD FOR VIA.
30-5=25 YEARS OLD FOR KYLE.
PROOF:
30+1+25+1=57
31+26=57
57=57